A jet leaves at the average rate of 564 km/hr. A half hour later, another jet leaves at 744km/hr. How long will it be until the second jet overtakes the first?
Answer by venugopalramana(3286)

A jet leaves at the average rate of 564 km/hr. A half hour later, another jet leaves at 744km/hr. How long will it be until the second jet overtakes the first?
in 1 hr.first jet travels 564 km.
hence initial distance of seperation = d = 564
relative speed..travelling in same direction = 744 - 564 =180 km/hr.
time = d / s = 564 /180 = 3.13 hrs.